Integrate this question please and and and and kindly simplify it cosx / tanx Integrate cosx/tanx

Grade:12

1 Answers

Nandana
6 years ago
 
hi
  here,
   ∫cosx/tanx dx = ∫ cos2 x/ sinx dx
                           =∫1-sin2 x /sinx dx
                           =∫1/sinx dx – ∫ sinx dx
                           = ∫ cscx dx + cosx
                  
                ∫ cscx dx = ∫ (csc2x +cotx cscx)/ (cscx+cotx)dx
                               
                                     u =cscx+cotx
                                    du = -(cotxcscx+csc2 x) dx
                               = -∫1/u du
                               = -lnu = ln(1/u) = ln(1/(cscx+cotx))
           
       ∫cosx/tanx dx = ln(1/(cscx+cotx)) +cosx +C is the answer !